OpenStudy (anonymous):

A city’s population is declining continuously at a rate of 6.1% each year. If the population of the city is currently 435,410, what will the population of the city be in 8 years? 222,930 263,162 266,945 267,278

OpenStudy (anonymous):

The easiest way to decrease something by 6.1 % (or 0.061) is to multiply it by (1 - 0.061) ...that is multiply it by 0.939 Do that multiplication 8 times.

OpenStudy (anonymous):

At the end of the first year the population will be: 435,410 * (1.0 - 0.061) = 435,410 * 0.939 At the end of the second year the population will be 435,410 * 0.939 * 0.939 At the end of the eighth year the population will be 4354100×0.9398 435410×0.9398=263161.9752 which rounds to 263,162